a lot has trees covering 33 1/3% of the area. if the trees cover 8166.7 ft.2 what is the total area to the nearest whole square foot.
Answer by stanbon(75887)

a lot has trees covering 33 1/3% of the area. if the trees cover 8166.7 ft.2 what is the total area to the nearest whole square foot.
---
Let total area = "x":
33 1/3% * x = 8166.7 ft^2
---
x = 8166.7/(33 1/3%)
----
x = (8166.7*100) / (33 1/3) = 24,500.1 sq. ft.
=========================
cheers,
Stan H.